### v2

# Three-dimensional chiral active Ornstein-Uhlenbeck model for helical motion of microorganisms

Abstract
Active movement is essential for the survival of microorganisms like bacteria, algae and unicellular parasites. In three dimensions, both swimming and gliding microorganisms often exhibit helical trajectories. Here we introduce a general stochastic dynamics model for these cases, namely a chiral active particle with an Ornstein-Uhlenbeck process for torque, which implies a finite correlation time for internal noise. We present an analytical solution which is in very good agreement with computer simulations. We then show that for this type of internal noise, chirality and rotation increase the persistence of motion and results in helical trajectories that have a larger long-time mean squared displacement than straight trajectories at the same propulsion speed. We finally provide an experimental example by analyzing imaging data for malaria parasites that glide through hydrogels on helical trajectories.
The survival of microbes like bacteria or algae is tightly connected to their ability to actively move, which is essential to seek out more favorable conditions, e.g. places which offer more nutrients or sunlight for photosynthesis [1, 2]. Although sometimes movement is collective, e.g. in biofilms or during swarming, at the heart of all migration processes is always the capability of single microbes to internally generate forces and torques [3, 4, 5]. Migration over large distances is also essential for unicellular eukaryotes that have specialized in infecting mammalian hosts, such as the causative agents of the diseases malaria or toxoplasmosis [7]. Interestingly, in three dimensions many microorganisms exhibit helical trajectories [8, 9]. This includes species of swimming bacteria [10, 11], swimming algae [12, 13, 14] and gliding parasites [15, 16, 17]. In Fig. 1, we show the helical trajectories of malaria parasites gliding through synthetic hydrogels (see Appendix A for experimental details).
Since the pioneering work of Berg and Purcell, it is well accepted that one of the main challenges for moving microorganisms is to counter the effects of stochastic noise [18, 19]. For example, the bacterium E. coli uses a run-and-tumble strategy to move up and down chemotactic gradients; here the tumble periods are believed to be required to reset the system because it tends to loose orientation during the run periods due to rotational Brownian noise [20]. The interplay of self-propulsion and external noise can be analyzed by the theory of active Brownian particles, which combines a propulsion force with stochastic noise for translation and rotation [21, 22, 23, 24]. The active Brownian particle model has been used and extended in various contexts [25, 26, 27, 28, 29, 30], including adding torques to obtain circle swimmers [31, 32, 33, 34, 35, 36, 37] and studying the influence of time-correlated noise [38, 39, 40, 41, 42].
Figure 1: Reconstructed trajectories of malaria parasites gliding through synthetic hydrogels. Because this environment is nearly isotropic, the right-handed helical trajectories persist for long times.
However, most of these studies considered 2D cases, whereas helical motion of chiral active particles occurs in 3D [28]. If chiral active motion in 3D was analyzed theoretically, then mostly in the context of swimming. An early work on asymmetric swimmers in 3D considered the analogy to polymer models to extract new power laws for effective diffusion [43]. Previous work on sperm swimming considered stochasticity on the level of curvature and torsion and showed that helical trajectories are useful search strategies for chemotaxis in noisy environments [44, 45]. This line of work also considered colored noise in the form of a power spectrum [45]. One study of chiral motion in 3D started from the full mobility tensor for an arbitrarily shaped particle and showed that helical trajectories are the most likely outcome [46]. Similarly, a chiral active Brownian particle model has been used to describe the helical motion of colonial choanoflagellates and to show that purely stochastic propulsion can result in effective dispersion [47]. Very recently, it has been shown in a deterministic model for sperm swimming that an asymmetric beat of the flagellum leads to helical trajectories with high persistence [48]. Collectively, this body of work demonstrates that helical trajectories can have evolutionary advantages for microorganisms moving in noisy environments.
Noise does not only arise from the interaction of the microorganisms with their environment, but also from the internal force generating processes, which have a correlation time on the same scale as the movement that they generate. To address this aspect of the system, Ornstein-Uhlenbeck (OU) processes have been used, usually replacing the body-fixed constant velocity with a noisy velocity performing an OU-process around a body-fixed average [49]. This naturally introduces a time scale to the noise and in the context of 2D swimmers in recent years has attracted substantial attention [50, 51, 52, 53, 54, 55, 56, 57]. However, the approach of using an OU-process has not yet been extended to chiral particles in 3D.
Here we introduce a three-dimensional model which represents the noise in the generation of torque as an OU-process, similarly as suggested previously for 2D [58]. This introduces a finite correlation time, reflecting transient but slower additional processes in the torque-generation mechanism, in contrast to uncorrelated (white) Brownian noise. For example, for gliding malaria parasites this could be some reorientation in the self-organized flow field of the adhesins on the parasite’s surface [17]. In the overdamped limit, where drag balances all forces and torques, the torque can be replaced by the resulting angular velocity. As we demonstrate here, this 3D-model for chiral active particles can be treated analytically by suitably truncating a hierarchy of equations. We derive equations for the effective correlation time, the mean position and the mean squared displacement (MSD). Our main finding is that, in 3D, chirality and hence rotation can lead to enhanced effective persistence compared to non-rotating particles by an integrative effect of stochastic noise; a stabilization that can even lead to helical trajectories becoming ”straighter than a straight line”, i.e. allowing for larger long-time MSD compared to a particle moving with the same speed on a straight trajectory without rotation. This suggests that helical trajectories are favored for microorganisms that have to quickly move large distances through their environment. Finally we compare our model to experimental data from malaria parasites, demonstrating that it can describe the MSD obtained from reconstructions of the imaged trajectories as shown in Fig. 1.
Model. To model the intrinsic rotational noise we consider an active particle that is moving with a body-fixed constant translational velocity $\mathbf{V}_{0}^{\text{body}}$. Its rotational velocity performs an OU-process around the body-fixed average, $\mathbf{\Omega}_{0}^{\text{body}}$. In the lab frame we use two vectors to track the orientation of the particle. $\mathbf{n}_{1}$ is the direction of the mean angular velocity $\mathbf{\Omega}_{0}$ and $\mathbf{n}_{2}\perp\mathbf{n}_{1}$ is defined with the angle $\alpha$ between $\mathbf{\Omega}_{0}$ and $\mathbf{V}_{0}$ (see Appendix Fig. A1 for a schematics):
$\mathbf{\Omega}_{0}=\Omega_{0}\mathbf{n}_{1}\ ,$ $\mathbf{V}_{0}=|\mathbf{V}_{0}|\left(\mathbf{n}_{1}\cos{\alpha}+\mathbf{n}_{2}\sin{\alpha}\right)\,.$ (1)
For simplicity, we set $|\mathbf{V}_{0}|=1$ in the following. In the lab frame, the equations of motion are
$\text{d}\mathbf{\Omega}\hskip 3.0pt=$ $\ -k\left(\mathbf{\Omega}-\Omega_{0}\mathbf{n}_{1}\right)\text{d}t+h\text{d}\mathbf{\Lambda}$ (2)
$\text{d}\mathbf{n}_{1}=$ $\ (\mathbf{\Omega}\times\mathbf{n}_{1})\text{d}t$ (3)
$\text{d}\mathbf{n}_{2}=$ $\ (\mathbf{\Omega}\times\mathbf{n}_{2})\text{d}t$ (4)
$\text{d}\mathbf{r}\hskip 6.0pt=$ $\ \left(\cos{(\alpha)}\mathbf{n}_{1}+\sin{(\alpha)}\mathbf{n}_{2}\right)\text{d}t\,.$ (5)
Here, $k$ is the potential strength and $h$ the noise amplitude of the OU-process. $\text{d}\mathbf{\Lambda}$ is a 3D standard Wiener process. Note that noise is not multiplicative in the lab frame. Focusing on the intrinsic noise for simplicity, we disregard external noise (such as Brownian translational noise) or intrinsic noise in the translational velocity.
Rotation. The rotational part described by Eq. 2-3 is decoupled from the rest and can be solved first. The dynamical equations for the expectation values $\left\langle\mathbf{\Omega}\right\rangle$ and $\left\langle\mathbf{n}_{1}\right\rangle$ constitute an infinite hierarchy of expectation values of cross products of these two quantities, the first four being
$\text{d}\left\langle\mathbf{\Omega}\right\rangle=$ $\ -k\left(\left\langle\mathbf{\Omega}\right\rangle-\Omega_{0}\left\langle\mathbf{n}_{1}\right\rangle\right)\text{d}t$ (6)
$\text{d}\left\langle\mathbf{n}_{1}\right\rangle=$ $\ \left\langle\mathbf{\Omega}\times\mathbf{n}_{1}\right\rangle\text{d}t$ (7)
$\text{d}\left\langle\mathbf{\Omega}\times\mathbf{n}_{1}\right\rangle=$ $\ -k\left\langle\mathbf{\Omega}\times\mathbf{n}_{1}\right\rangle\text{d}t+\left\langle\mathbf{\Omega}\times(\mathbf{\Omega}\times\mathbf{n}_{1})\right\rangle\text{d}t$ (8)
and
$\text{d}\left\langle\mathbf{\Omega}\times(\mathbf{\Omega}\times\mathbf{n}_{1})\right\rangle=$ $\ -2k\left\langle\mathbf{\Omega}\times(\mathbf{\Omega}\times\mathbf{n}_{1})\right\rangle\text{d}t$
$\ +\left\langle\mathbf{\Omega}\times\left(\mathbf{\Omega}\times(\mathbf{\Omega}\times\mathbf{n}_{1})\right)\right\rangle\text{d}t$
$\ +k\Omega_{0}\left\langle\mathbf{n}_{1}\times(\mathbf{\Omega}\times\mathbf{n}_{1})\right\rangle\text{d}t$
$\ +h^{2}\left\langle\text{d}\mathbf{\Lambda}\times(\text{d}\mathbf{\Lambda}\times\mathbf{n}_{1})\right\rangle\ .$
To obtain an analytical solution, we apply moment closure by approximating the second and third term in this equation. The truncation cannot be performed earlier due to the necessity of retaining terms up to second order in $\Omega$ to properly account for the effect of noise. First, assuming $\left\langle\mathbf{\Omega}^{2}A\right\rangle\approx\Omega_{0}^{2}\left\langle A\right\rangle$, we get for the second term $\left\langle\mathbf{\Omega}\times\left(\mathbf{\Omega}\times(\mathbf{\Omega}\times\mathbf{n}_{1})\right)\right\rangle\approx-\Omega_{0}^{2}\left\langle\mathbf{\Omega}\times\mathbf{n}_{1}\right\rangle$. For the third term, upon replacing $\mathbf{\Omega}$ with $\mathbf{\Omega}^{\perp}$, its component perpendicular to $\mathbf{n}_{1}$, the relevant contribution for the cross product, and applying similar logic as before but with $|\mathbf{n}_{1}|=1$, we get $\left\langle\mathbf{n}_{1}\times(\mathbf{\Omega}\times\mathbf{n}_{1})\right\rangle=\left\langle\mathbf{\Omega}^{\perp}\right\rangle=\left\langle\mathbf{\Omega}-(\mathbf{\Omega}\cdot\mathbf{n}_{1})\mathbf{n}_{1}\right\rangle\approx\left\langle\mathbf{\Omega}\right\rangle-\Omega_{0}\left\langle\mathbf{n}_{1}\right\rangle$. For the last approximation, we assumed that the variance of the OU-process is small, such that $\mathbf{\Omega}$ stays close to its average $\Omega_{0}\mathbf{n}_{1}$. The noise term can be explicitly computed, and we can close the hierarchy by rewriting its fourth equation as
$\text{d}\left\langle\mathbf{\Omega}\times(\mathbf{\Omega}\times\mathbf{n}_{1})\right\rangle=-2k\left\langle\mathbf{\Omega}\times(\mathbf{\Omega}\times\mathbf{n}_{1})\right\rangle\text{d}t$ (9)
$\ -\Omega_{0}^{2}\left\langle\mathbf{\Omega}\times\mathbf{n}_{1}\right\rangle\text{d}t+k\Omega_{0}\left(\left\langle\mathbf{\Omega}\right\rangle-\Omega_{0}\left\langle\mathbf{n}_{1}\right\rangle\right)\text{d}t-2h^{2}\left\langle\mathbf{n}_{1}\right\rangle\text{d}t.$
By rotational symmetry, only the component singled out by the initially parallel $\mathbf{\Omega}$ and $\mathbf{n}_{1}$ axes is relevant (as the rotational problem is independent of $\mathbf{n}_{2}$), and the other two components of each vector vanish upon averaging (we choose this direction to be $z$). Hence, the truncated system defines a four-dimensional, linear, homogeneous ordinary differential equation problem, which we can analyze by its eigenvalues (see Appendix B for the truncation in the case of vanishing $\Omega_{0}$).
The mode relevant for the long time behavior can be identified as the unique mode with real eigenvalue and parallel $\left\langle\mathbf{\Omega}\right\rangle$ and $\left\langle\mathbf{n}_{1}\right\rangle$, which describes the decorrelation of $\left\langle\mathbf{\Omega}\right\rangle$ from its initial orientation. The other eigenvalues describe the unstable state where $\mathbf{\Omega}$, $\mathbf{n}_{1}$ are anti-parallel, and oscillatory states, all of which decay more quickly. The relevant eigenvalue can be computed exactly, but as a solution of a fourth-order polynomial, it is cumbersome. Expanding for small $k/\Omega_{0}$, i.e. assuming the rotation is faster then the timescale on which the OUP returns to its average, we get the approximation
$\lambda=-\frac{h^{2}+\Omega_{0}^{2}k+k^{3}-\sqrt{k^{2}\left(\Omega_{0}^{2}+k^{2}\right)^{2}-h^{4}}}{\Omega_{0}^{2}+k^{2}}\ \,$ (10)
which describes the decay as $\left\langle\mathbf{\Omega}\right\rangle=(0,0,\Omega_{0})\exp{(\lambda t)}$ and $\left\langle\mathbf{n}_{1}\right\rangle=(0,0,1)\exp{(\lambda t)}$. The more negative $\lambda$, the faster $\mathbf{\Omega}$ and $\mathbf{n}_{1}$ decorrelate from their initial orientation. For fixed noise amplitude $h$, both decreasing the strength $k$ of the OUP potential and decreasing angular speed $\Omega_{0}$ lead to faster decorrelation, suggesting that the rotation has a stabilizing effect. In the limits of vanishing noise, diverging potential strength or diverging angular speed, the time scale of decorrelation diverges. In the limit of small noise amplitude, $\lambda$ converges to the power spectrum of the OU-process evaluated at the angular speed, consistent with a derivation starting from power spectra ([45], see Supplemental Material).
Figure 2: a: Time course of different moments obtained from numerical simulation (Eq. 2-5) in comparison with numerical solution of the truncated system (Eq. 6-9, gray dashed lines) and analytical approximation predicting exponential decay with eigenvalue $\lambda$ (Eq. 10, yellow dotted lines). Parameter values: potential strength $k=$ 0.2, noise amplitude $h=$ 0.3, angular speed $\Omega_{0}=$ 1, angle $\alpha=\pi/6$. b: Same as a, but now for $k=$ 2, $h=$ 0.1 and $\Omega_{0}=$ 2, i.e. much reduced noise and faster turning. Here, the agreement between simulations and theory is even better. c: Simulated trajectories at parameters from a. d: Simulated trajectories at parameters from b. The reduced noise leads to more regular trajectories. See also Supp. Movies 1+2.
To validate our approximations, we compared the solutions against numerical simulations of the initial model (Eq. 2-5, implemented in JAX [59] using standard solvers for stochastic differential equations, set up in diffrax [60]). In Fig. 2, different expectation values obtained from averaging 20.000 numerical simulations are compared with (i) the numerically solved truncated ODE system (Eq. 6-9, dashed gray) and (ii) the analytical exponential decay given by the dominant eigenvalue $\lambda$. As shown in Fig. 2a, larger noise yielding faster decorrelation produces larger values of the higher order expectation values. We find that the truncations are qualitatively correct (see also Supplemental Figure S1), while quantitative differences are visible – the numerical solution of the truncated system shows some additional oscillations. The exponential decay by $\lambda$ is too fast here, which signifies that during the relatively rapid decay, additional modes are relevant. For lower noise, in Fig. 2b we see excellent agreement between numerical simulation, numerical solution of the truncated system, and the exponential decay given by $\lambda$ from Eq. 10. The latter decay completely dictates the persistence in the resulting motion, as also apparent in the resulting trajectories illustrated in Fig. 2c+d. Generally, lower decorrelation can be reached by lower noise amplitude $h$, stronger Ornstein-Uhlenbeck potential $k$, or higher angular speed $\Omega_{0}$.
Translation. For the analytical treatment of the translational part, we assume that initially $\mathbf{\Omega}=\Omega_{0}\mathbf{n}_{1}$ is in $z$ direction. The solution of the rotational part then allows to solve Eq. 5 for the motion in $z$, $\text{d}\left\langle z\right\rangle=\ \cos{(\alpha)}\left\langle\mathbf{n}_{1}\right\rangle$. To obtain the MSD and the remaining coordinate, we need an expression for $\left\langle\mathbf{n}_{2}\right\rangle$. By construction, $\mathbf{n}_{2}\perp\mathbf{n}_{1}$, so $\mathbf{n}_{2}$ is rotating in the plane perpendicular to $\mathbf{n}_{1}$ with angular frequency $\Omega_{0}$, which is on average the $x$-$y$-plane. We assume that $\mathbf{n}_{2}$ initially points in $x$-direction. The decorrelation of $\mathbf{n}_{1}$ is also decorrelating the plane in which $\mathbf{n}_{2}$ rotates, but the latter additionally decorrelates within the plane by variations of the magnitude of the rotational velocity. Both effects are caused by $\mathbf{\Omega}$ deviating from $\mathbf{\Omega}_{0}$, the tilting of the plane by deviations perpendicular to $\mathbf{\Omega}_{0}$, and in-plane deviations by parallel components. Because of this additional effect, we obtain a two-fold faster decorrelation of $\left\langle\mathbf{n}_{2}\right\rangle$ compared to $\left\langle\mathbf{n}_{1}\right\rangle$ (see Appendix C for more details),
$\left\langle\mathbf{n}_{2}\right\rangle=\big{(}\cos{\Omega_{0}t},\sin{\Omega_{0}t},0\big{)}\,e^{2\lambda t}\ .$ (11)
The MSD can now be obtained (see Appendix D) by first computing it from the formal solution and inserting the solutions obtained for $\left\langle\mathbf{n}_{1}\right\rangle$ and $\left\langle\mathbf{n}_{2}\right\rangle$. The result is (except for the degenerate case $\Omega_{0}=0$ and $\alpha>0$):
$\begin{split}&\left\langle\mathbf{r}^{2}_{(t)}\right\rangle=\frac{2\cos^{2}(\alpha)\left(-\lambda t+e^{\lambda t}-1\right)}{\lambda^{2}}\\ &\ +\frac{2\sin^{2}(\alpha)}{\left(4\lambda^{2}+\Omega_{0}^{2}\right)^{2}}\Big{[}-4\lambda^{2}+\Omega_{0}^{2}-2\lambda\left(4\lambda^{2}+\Omega_{0}^{2}\right)t\\ &\ +\left(4\lambda^{2}-\Omega_{0}^{2}\right)e^{2\lambda t}\cos(\Omega_{0}t)+4\lambda\Omega_{0}e^{2\lambda t}\sin(\Omega_{0}t)\Big{]}\ .\end{split}$ (12)
Let us consider two limiting cases. First, for $\alpha=0$, corresponding to a particle rotating while traveling in average on a straight line, we obtain
$\left\langle\mathbf{r}_{(t)}^{2}\right\rangle=-\frac{2}{\lambda}t-\frac{2}{\lambda^{2}}\left(1-e^{\lambda t}\right)\ ,$ (13)
which recovers the case of an active Brownian particle (see Supplemental Material, note $\lambda<0$). Second, for general $\alpha$ in the limit of large $t$, we can approximate $\left\langle\mathbf{r}_{(t)}^{2}\right\rangle\approx 6D_{\infty}t$, where we obtain the effective diffusion constant describing the long time behavior as
$D_{\infty}=-\frac{\lambda}{3}\left(\frac{\cos^{2}(\alpha)}{\lambda^{2}}+\frac{2\sin^{2}(\alpha)}{\left(4\lambda^{2}+\Omega_{0}^{2}\right)}\right)\ .$ (14)
In the case $\lambda^{2}\ll\Omega_{0}^{2}$, meaning small noise leading to a decay time much longer than the rotation period, and $\alpha<\pi/2$, i.e. the particle not just circling, but having some average net movement, this reduces to $D_{\infty}\approx-\cos^{2}(\alpha)/(3\lambda)$.
Lastly, with Eq. 11 we compute $\left\langle x\right\rangle$ and $\left\langle y\right\rangle$ by integrating $\text{d}\left\langle x\right\rangle=\sin(\alpha)\left\langle\mathbf{n}_{2}\right\rangle_{x}\text{d}t$, allowing to obtain the expectation value of the trajectory,
$\left\langle\mathbf{r}_{(t)}\right\rangle=\begin{pmatrix}\sin(\alpha)\frac{e^{2\lambda t}(2\lambda\cos(\Omega_{0}t)+\Omega_{0}\sin(\Omega_{0}t))-2\lambda}{4\lambda^{2}+\Omega_{0}^{2}}\\ \sin(\alpha)\frac{e^{2\lambda t}(2\lambda\sin(\Omega_{0}t)-\Omega_{0}\cos(\Omega_{0}t))+\Omega_{0}}{4\lambda^{2}+\Omega_{0}^{2}}\\ \frac{\cos{(\alpha)}}{\lambda}\left(e^{\lambda t}-1\right)\end{pmatrix}\ ,$ (15)
which is a logarithmic spiral on a radical surface, i.e. $z\propto\sqrt{r}$.
Figure 3: a: Mean distance traveled in $z$-direction (the initial orientation of the helical axis) for different $\Omega_{0}$ at $k=$ 2, $h=$ 0.1. Full lines show particles moving straight while turning ($\alpha$=0), dashed lines particles on helical trajectories ($\alpha$=$\pi/4$), which can be seen overtaking slower turning straight particles. Colored and black lines are theoretical and numerical, respectively, and in very good agreement. b: Mean squared displacement for the same parameters as shown in a, theoretical results from Eq. 12 in color. c: Effective long-time diffusion constant $D_{\infty}$, cf. Eq. 14, as a function of noise amplitude $h$ and angular speed $\Omega_{0}$. Black lines mark contours of constant $D_{\infty}$.
Results. The derived solutions show that increasing rotation $\Omega_{0}$ stabilizes the particle against its intrinsic noise. In Fig. 3a and b we plot the mean $z$-position and the MSD, respectively, which both increase with increasing angular speed. In both cases we find that the numerical simulations agree well with the analytical results, cf. the third component of Eq. 15 and Eq. 12. The plots show that if the particle travels on a helix (case $\alpha=\pi/4$, cf. dashed lines in Fig. 3a,b) with the same speed as a non-rotating particle traveling in a straight fashion, if it turns sufficiently fast (i.e. if the helix is sufficiently tightly wound) it travels further from the origin on average at long time scales. Therefore a helical trajectory can be ”straighter than a straight line”.
The long time behavior is described by the effective diffusion constant Eq. 14, which has a complicated dependence on $\Omega_{0}$, $k$ and $h$ through $\lambda$. In Fig. 3c we see that at constant OU potential strength $k$, $D_{\infty}$ increases with higher angular speed $\Omega_{0}$, as this suppresses deviations of the helical axis, different from what was found for chiral active Brownian particles without the OU-process [47] (see also Supplemental Material). This effect becomes more pronounced for higher noise amplitude $h$, i.e. at higher noise, the stabilizing effect of rotation is more pronounced. Increasing effective diffusion by introducing rotation or equivalently chirality is strikingly different from known examples. In 2D, chirality reduces long-time diffusion by enforcing circular turning [35, 61]. Similarly, a 3D active Brownian particle with external torque exhibits reduced long-time diffusion [28].
Figure 4: a: Mean squared displacement for $\alpha$ close to $\pi/2$, such that the particles are close to describing circles, with $k=$ 1, $h=$ 0.5, $\Omega_{0}=$ 2. Black dotted lines are averages from numerical simulations. b: Theoretical expectation value of trajectories (Eq. 15) for the two lower values of $\pi/2-\alpha$.
Figure 5: a: Log-log plot of the mean squared displacement extracted from observed malaria parasite trajectories shown in Fig. 1 (purple, dotted), with five percent percentiles (purple, shaded) and the fitted model (orange). The gray dashed line is a fitted power law. b: Deviation from fitted power law. The vertical dashed line marks one period of rotation as extracted from the fitted model. c: Trajectories simulated with parameters obtained from MSD fit resembling Fig. 1 (cf. Supp. Movie 3).
We can also study the short time behavior. Fig. 4a shows that at short times, the MSD of a helix grows slower, because it is curving back onto itself, depending on the pitch of the helix defined by the angle $\alpha$. For $\alpha=\pi/2$, the MSD shows strong oscillations, as the mean position, see Fig. 4b, describes a planar inward spiral due to the influence of noise that diverts it from the circle of a noise-free particle. For smaller $\alpha$, the spiral gets the 3D structure of a logarithmic spiral on a radical surface as found in Eq. 15, with both cases showing good agreement between the numerical and analytical results.
Finally, we can use our measured trajectories for malaria parasites in hydrogels to extract their MSD (Fig. 5a, averaged from 140 trajectories) and fit it with our model prediction of Eq. 12, similar to what has been done before in 2D-projections for choanoflagellate colonies [47]. In general, we find good agreement. Our theory successfully describes the first two extrema in the deviation of the MSD from a power law (Fig. 5b), corresponding to the first turn of the helix. Our theory also predicts some effects of second and third turns visible in the MSD deviation, which are not observable in the experimental data, most likely because the biological population has a distribution of helical pitches and radii such that the later turns cannot be resolved in the average. From the fitted model parameters (Appendix E), we can derive estimates for pitch and radius of the helical trajectories as 13.2 $\upmu$m and 2.8 $\upmu$m, respectively, well within the observed range [15]. We note in passing that a phenomenological fit of the period as the period in the MSD-deviation would yield incorrect results for the oscillation period. Trajectories simulated with the fitted parameters (Fig. 5c) visually resemble the observed trajectories from Fig. 1.
In summary, our results suggest that helical trajectories provide an advantage for swimming or gliding microorganisms with noisy force generation to effectively cover distance more quickly than when going straight. This result from a general stochastic theory for microorganisms with colored noise in their internal torque generation complements early insight about the potential evolutionary advantages of helical motion for swimming [44, 47, 48]. In the future, our model could guide the design of micro- and nanobots [62, 63], for example in medical applications where enhanced persistence of motion is required [64].
Acknowledgments: This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through Priority Programme 2332 (Projektnummer 492010213) and Collaborative Research Center 1129 (Projektnummer 240245660).
Data availability: The code used to simulate the model, as well as python implementations of the analytical solutions for comparison are available in the following repository: GITHUB URL UPON PUBLICATION
Appendix D - Derivation of MSD: The MSD can now be obtained by first computing it from the formal solution
$\mathbf{r}_{(t)}=\int_{0}^{t}\text{d}t\left(\cos{(\alpha)}\mathbf{n}_{1}+\sin{(\alpha)}\mathbf{n}_{2})\right)\ ,$ (24)
which yields
$\begin{split}\left\langle\mathbf{r}_{(t)}^{2}\right\rangle=\int_{0}^{t}\text{d}s_{1}&\int_{0}^{t}\text{d}s_{2}\Big{(}\cos^{2}{(\alpha)}\left\langle\mathbf{n}_{1}(s_{1})\cdot\mathbf{n}_{1}(s_{2})\right\rangle\\ &+\sin^{2}{(\alpha)}\left\langle\mathbf{n}_{2}(s_{1})\cdot\mathbf{n}_{2}(s_{2})\right\rangle\Big{)}\ ,\end{split}$ (25)
where the mixed terms vanish, as due to their perpendicularity and the rotational symmetry the expectation value of their scalar products has to be zero even if evaluated at different times. The remaining correlation functions can be directly obtained from the solutions obtained for $\left\langle\mathbf{n}_{1}\right\rangle$ and $\left\langle\mathbf{n}_{2}\right\rangle$ as
$\left\langle\mathbf{n}_{1}(s_{1})\cdot\mathbf{n}_{1}(s_{2})\right\rangle=$ $\ e^{\lambda|s_{1}-s_{2}|}\ ,$ (26)
$\left\langle\mathbf{n}_{2}(s_{1})\cdot\mathbf{n}_{2}(s_{2})\right\rangle=$ $\ \cos{(\Omega_{0}|s_{1}-s_{2}|)}e^{2\lambda|s_{1}-s_{2}|}\ ,$ (27)
such that we finally compute the MSD by simple integration.
Appendix E - Fitted Parameters: Experimentally observed sporozoites do not move at a constant speed, and even their average speed can vary between different parasites by a factor of 3. For this analysis, we resampled the trajectories assuming a constant speed of 1 $\upmu$m/s, which gives more structures to the predicted MSD than observed experimentally. The fit of the MSD in Fig. 5 results in the parameters given in Table 1.
Table 1: Results for fitting the model to the MSD computed from experimentally observed and resampled trajectories of malaria parasites.
Parameter $\Omega_{0}$ $\alpha$ $\theta$ $h$
Fit Result 0.285 1/s 0.926 1.078 1/s 0.154 1/s3/2